# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 1.4 Derandomization for short regular ROBPs
There exists an algorithm that takes a regular ROBP $f$ of length $n$, width $w$ and alphabet $\{0,1\}^{s}$ and a parameter $\varepsilon>0$ as input, and outputs a value that approximates $\mathbb{E}[f]$ within error $\varepsilon$. The algorithm runs in $O(s+\log n(\log\log n+\sqrt{\log(w/\varepsilon)})+\log(w/\varepsilon))$ bits of space.
Notice that when $n\leq 2^{O(\log^{1/2}w)}$ and $\varepsilon=1/\operatorname{\mathrm{poly}}w$ our algorithm achieves a space complexity of $O(s+\log w)$, which is optimal up to a constant factor, i.e. in $\mathbf{L}$ exactly. When requiring the derandomization to be in $\mathbb{L}$, the only previous method that can work is actually the Nisan-Zuckerman generator [NZ96] and also its variation in [Arm98]. Our result significantly improves the length and error of the regular ROBPs that can be derandomized in this case. On the other hand, if comparing with [AKM+20, CHL+23, CL24], then our result not only improves the space complexity to optimal, but also improves the time complexity from super-polynomial to polynomial. In fact, we note that even if $n=\operatorname{\mathrm{poly}}\log w$, the derandomizations of [AKM+20, CHL+23, CL24] are still in super-polynomial time.
One may want to further improve the randomness dependence on $\varepsilon$ by first using our theorem with error $1/\operatorname{\mathrm{poly}}w$ and then calling the Richardson iteration method of [AKM+20]. This can be done, but our main purpose here is to get a derandomization in $\mathbb{L}$ exactly. Note that if applying that additional operation, then the algorithm is not in $\mathbb{L}$ and also not in polynomial time.
@@ 1.5 Technical Overview
Our general method is inspired first by an observation about the error reduction technique introduced in [CDR+21, PV21]. Recall that in [CDR+21, PV21], a given ROBP $\{A_{i}\}_{i=1}^{n}\in\mathbb{R}^{w\times w}$, the product $A=\Pi_{i=1}^{n}A_{i}$ can be approximated by a weighted summation of a sequence of iterated matrix multiplications,
$\left\lVert A-\sum_{i\in[K]}\sigma_{i}\cdot B_{0,n_{i,1}}B_{n_{i,1},n_{i,2}}\ldots B_{n_{i,k-1},n_{i,k}}\right\rVert\leq\varepsilon_{0}^{k}\cdot(n+1).$
[…]
An observation is that $B_{0,n_{i,1}}B_{n_{i,1},n_{i,2}}\ldots B_{n_{i,k-1},n_{i,k}}$ can be viewed as a shorter ROBP with a large alphabet, while the expectation of the original branching program can be approximated as the sum of weighted expectations of shorter branching programs. So an initial thought is that if one can somehow recursively apply this procedure, to further reduce the length of the ROBP, until the length becomes a constant, then by using true randomness for constant length ROBPs, one can attain a WPRG. Of course, there are problems with this initial thought: first, the alphabet size may increase quickly; second, one needs to make sure the length can indeed be shorter each time we apply the reduction; third, the overall weight needs to be bounded. We will exhibit how to resolve these problems respectively for our results.
To describe our constructions in detail, we introduce a natural concept, the weighted pseudorandom reduction, which is defined as follows.
@@
Let $\mathcal{F}_{0}$ be a class of functions ($\left(\{0,1\}^{s_{0}}\right)^{n_{0}}\rightarrow\mathbb{R}$. Let $\mathcal{F}_{simp}$ be a class of functions ($\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\rightarrow\mathbb{R}$. A $(d,K,\varepsilon)$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{simp}$ is a tuple $(\mathsf{R},w)$, in which $\mathsf{R}:\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\times\{0,1\}^{d}\rightarrow\left(\{0,1\}^{s_{0}}\right)^{n_{0}}$ and $w:\{0,1\}^{d}\rightarrow\mathbb{R}$. Furthermore, for every $f\in\mathcal{F}_{0}$, we have:
@@
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d}}\sum_{i\in\{0,1\}^{d}}w(i)\mathbb{E}_{U^{\prime}}[f(\mathsf{R}(U^{\prime},i))]\right|\leq\varepsilon,$
$\forall i,|w(i)|\leq K,$
$\forall i,f(\mathsf{R}(\cdot,i))\in\mathcal{F}_{simp}.$
A weighted pseudorandom reduction can be viewed as a ‘partial assignment’ for the seed of a weighted pseudorandom generator. Instead of directly approximating $\mathbb{E}[f]$ with $\sum_{r}\frac{1}{2^{s}}w(r)\cdot f(r)$, we consider approximating $\mathbb{E}[f]$ with $\sum_{i}\frac{1}{2^{d}}w(i)\cdot\mathbb{E}_{U_{s_{1}}}[f(\mathsf{R}_{i}(U_{s_{1}}))]$. So for each $i$, we only need to further approximate $\mathbb{E}[f(\mathsf{R}_{i}(U_{s_{1}}))]$ with another weighted pseudorandom reduction. Despite its resemblance to error reduction frameworks, the weighted pseudorandom reduction is not designed for recursively reducing the error. Instead, the core idea is to approximate a complex ROBP by a weighted sum of simpler ROBPs (e.g. shorter or smaller alphabet), and recursively repeat this process until the ROBPs are simple enough to be handled easily.
WPRGs for standard-order ROBPs
To describe our construction, we start by reviewing Armoni’s generator in view of weighted pseudorandom reductions. For simplicity, we consider $\varepsilon\geq 1/w^{3}$. Armoni’s generator can be viewed as a recursive weighted pseudorandom reduction procedure. It has $k$ levels of recursions. Let $f$ be the original ROBP. At level $i$ of the recursion, assume the current ROBP is $f_{i}$ of length $n$ and alphabet $\{0,1\}^{C\log w}$, where $C>0$ is a constant. We prepare a source $X$ of $O(\log(nw/\varepsilon))$ bits and extracts it using multiple independent seeds $Y_{1},\ldots,Y_{n}$, each of $O(\log(n/\varepsilon))$ bits. This outputs $n$ chunks of $(C\log w)$-bit symbols. And the number of random bits used is reduced from $Cn\log w$ to $O(n\log(n/\varepsilon))$. Notice that this is also a one-level NZ generator. If we further fix the source $X$, the composition of $f_{i}$ and the one-level NZ generator can be viewed as a new ROBP that operates on $Y_{1},\ldots,Y_{n}$, which can be further regarded as a ROBP of length $m=n\cdot\frac{O(\log(n/\varepsilon))}{C\log w}$ and alphabet $\{0,1\}^{C\log w}$. This perspective allows one to recursively apply the process. At each level $i$ of recursion, one can reduce a ROBP of length $n_{i}$ and alphabet $\{0,1\}^{C\log w}$ to a ROBP of length $n_{i+1}=n_{i}\cdot\frac{O(\log(n/\varepsilon))}{C\log w}$ and the alphabet $\{0,1\}^{C\log w}$. The cost of this level is that we use $O(\log(nw/\varepsilon))$ fresh random bits as a source to extract. By $k$ levels of recursion, Armoni’s generator achieves a seed length of $O\left(k\cdot\log(nw/\varepsilon)+n\cdot\left(\frac{\log(n/\varepsilon)}{\log w}\right)^{k}\right)$ for any integer $k\geq 1$. However, when $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$, the multiplicative factor $\frac{O(\log(n/\varepsilon))}{C\log w}$ deteriorates to a constant. Under such cases, one can set $k$ to be $\log n$ to attain a short seed. But then the seed length is no better than just applying Nisan’s generator.
The first ingredient in our construction is the length reduction, which can accelerate the length-decreasing rate in the above procedure. For simplicity, we first focus on $n=\operatorname{\mathrm{poly}}\log w$. We will handle larger $n$ later. We show how to make $n_{i+1}=n^{1/3}_{i}$ and how to attain an advantage in seed length based on this property. By a standard method from [CDR+21, PV21] based on Richardson iteration, one can transform a PRG with a moderate error $\tau$ into a WPRG with a tiny error $\varepsilon$. For level $i$, we set the moderate error $\tau=2^{-\frac{C\log w}{n^{1/c}}}$ for some constants $C,c>0$. The WPRG constructed in this manner has the form:
@@
$w(j)=\sigma_{j}\cdot K.$
Here $m=O\left(\frac{\log(n/\varepsilon)}{\log(1/\tau)}\right)$, $j$ is chosen from a predefined set $[K]$, $K=n^{O(m)}$, $0\leq n_{j,1}<n_{j,2}<\ldots<n_{j,m}\leq n$ are breakpoints, and $\sigma_{j}\in\{-1,0,1\}$. All these parameters are from a certain preconditioned Richardson iteration. The PRG $\mathsf{PRG}(\cdot)$ that we use here, is the NZ generator with $c$ levels of recursion, maintaining a seed length of $O(\log w)$. The output sequence is a concatenation of $m$ chunks of independent NZ generator outputs, each truncated to a length of $n_{j,t}-n_{j,t-1}$ bits. To repeat the process, we convert the WPRG into a weighted pseudorandom reduction by fixing $j$ and allowing $x_{1},\ldots,x_{m}$ to remain free. Namely, we define $(\mathsf{R},w)$, with $\mathsf{R}_{j}(x)=\mathsf{PRG}(x_{1})_{n_{j,1}},\mathsf{PRG}(x_{2})_{n_{j,2}-n_{j,1}},\ldots,\mathsf{PRG}(x_{m})_{n_{j,m}-n_{j,m-1}}$ and $w(j)=\sigma_{j}\cdot K$. The reduced ROBP can be viewed as running on $m$ large symbols, each corresponding to a seed of an NZ generator $\mathsf{PRG}(\cdot)$. The key point is that by the standard calculation in [CDR+21, PV21], $m=O\left(\frac{\log(n/\varepsilon)}{\log(1/\tau)}\right)=n^{1/c}$, if $\varepsilon=1/\operatorname{\mathrm{poly}}w$ and we take the constant $C$ in $\tau$ to be large enough. Hence if we take $c=3$, $n=n_{i}$, then $n_{i+1}=n_{i}^{1/3}$.
In principle, this process can be repeated until $n_{i}$ is reduced to a constant, requiring $l=\log\log n=\log\log\log w$ levels of recursion. However, when we apply the NZ generator for a ROBP with a large alphabet $\{0,1\}^{s}$, it requires a seed of $\Theta(s+\log w)$ bits, the constant hidden in big-$\Theta$ can at least double the bit-length of the alphabet. Hence such a direct repetition may lead to a final alphabet of $\log w\cdot 2^{l}=O(\log w\log\log w)$ bits, which loses the advantage in seed length.
To solve this problem, we introduce another reduction, the alphabet reduction, which can control the size of the alphabet using a one-level NZ generator. Assume we have a ROBP $f$ with length $m$ and alphabet log-size being some large $s$. Recall that a one-level NZ generator, on an input source $X$ of $O(s+\log w)$ bits, and a sequence of independent seeds $Y_{1},\ldots,Y_{m}$, each of $d=O(\log(n/\varepsilon))$ bits, can output $m$ symbols each of length $s$. So specifically, we can define $(\mathsf{R},w)$, where $\mathsf{R}_{x}(y_{1},y_{2},\ldots,y_{m})=\mathrm{E\textsc{xt}}(x,y_{1}),\mathrm{E\textsc{xt}}(x,y_{2}),\ldots,\mathrm{E\textsc{xt}}(x,y_{m})$ and $w(x)\equiv 1$. Hence by fixing $X=x$, then $f(\mathsf{R}_{x})(\cdot)$ is a ROBP with length $m$ and alphabet log-size $d$ which can be a small constant fraction times $|x|$.
Now we describe our whole recursive reduction that alternates between length reductions and alphabet reductions. Let $\varepsilon=1/\operatorname{\mathrm{poly}}(w)$. We will handle large $n$ later, but now assume at some step we have already obtained ROBPs of length $n_{i}\leq\log^{2}w$, alphabet log-size $s_{i}=C_{1}\log w$, and width $w$. Next, we first apply the length reduction, which gives us ROBPs of length $n_{i+1}=n^{1/3}_{i}$, width remaining the same, and alphabet log-size $O(s_{i})=C\log w$ for some large constant $C$. The weight of this reduction is $K=n_{i}^{O(n_{i+1})}\leq 2^{O(n_{i+1}\log n_{i})}$, and we need $\log K$ random bits to choose an addend. Then we apply the alphabet reduction, which gives us ROBPs of the same length $n_{i+1}$, width $w$, but alphabet log-size $s_{i+1}=O(\log(n_{i+1}/\varepsilon))=C_{1}\log w=s_{i}$. This reduction requires $O(\log w)$ bits of randomness as a source for the extractor. Combining these two steps, we reduce $n_{i}$ to $n_{i+1}$ and maintain the same width and alphabet. The overall cost is $O(\log w)$ bits of randomness and weight $W\leq 2^{O(n_{i+1}\log n_{i})}$. One can repeat this process until $n$ is small enough, and this gives our new WPRG. A detailed calculation can deduce that the whole process needs randomness $l\cdot O(\log w)=O(\log w\log\log\log w)$. The overall weight $W$ is the product of weights from each level, i.e. $W=2^{\sum_{i}O(n_{i+1}\log n_{i})}\ll\operatorname{\mathrm{poly}}w$. Also, notice that one can let $\varepsilon$ be our target error times $1/W$ to bound the overall error.
As mentioned before, the above approach works for a ROBP of length $n=\log^{2}w$. To extend it to a large length, we use an extra length reduction based on Armoni’s PRG and the standard error reduction [CDR+21, PV21], to reduce the original ROBP to a ROBP that has length $<<O(\log^{2}w)$, which fits into the regime we have already settled.
Furthermore, to achieve a smaller target error $\varepsilon\ll 1/\operatorname{\mathrm{poly}}(w)$, we slightly modify the sampler technique given by Hoza [Hoz21], to work on our above construction. We mention that this operation can also be viewed as an extra level of weighted pseudorandom reduction. This ultimately gives the WPRG in Theorem 1.5.
We start by reviewing the recent WPRG [CHL+23] in a weighted pseudorandom reduction view. Their WPRG combines the INW generator and a new matrix iteration. To fool a unbounded-width permutation ROBP of length $n$ and alphabet $\{0,1\}^{s}$ within error $\varepsilon$, they first prepare a INW generator $\mathsf{PRG}:\{0,1\}^{d}\to\left(\{0,1\}^{s}\right)^{n}$ with moderate sv-error [^5] See Definition 4.5. $\tau=2^{O(\log\log n\sqrt{\log(1/\varepsilon)})}$. Then they construct the $\varepsilon$-WPRG $(G,w)$ in the following form:
$w(j)=\sigma_{j}\cdot K.$
@@
Here $m=O\left(\log n\cdot\frac{\log(\log n/\varepsilon)}{\log(1/\tau)}\right)$, $j$ is choosed from set $[K]$, $K=2^{O(m)}$, $0\leq n_{j,1}<n_{j,2}<\ldots<n_{j,m}\leq n$ are the breakpoints, and $\sigma_{j}\in\{-1,0,1\}$. All these parameters are from their new matrix iteration. The seed length of such INW generator is $d=s+O\left(\log n\cdot(\log\log n+\log(1/\tau))\right)$. Denote $(G,w)$ with the above parameters as $\mathsf{R}^{(0)}$. Viewing $\mathsf{R}^{(0)}$ as a weighted pseudorandom reduction, it reduces the original ROBP to ROBPs of length $m$ and alphabet $\{0,1\}^{d}$. Finally, they use an $\varepsilon/K$-INW generator to fool these reduced permutation ROBPs, which cost $d+O\left(\log m\cdot(\log\log m+\log(K/\varepsilon)\right)$ bits of randomness. This step introduces some double-logarithmic factors.

### v3

@@ 1.4 Derandomization for short-wide regular ROBPs
There exists an algorithm that for any input regular ROBP $f$ of width $w$, length $n=2^{O(\log^{1/2}w)}$, alphabet $\{0,1\}^{s}$, and for any input parameter $\varepsilon\geq 1/\operatorname{\mathrm{poly}}w$, outputs an approximation for $\mathbb{E}[f]$ with addtive error $\varepsilon$. The algorithm uses $O(s+\log w)$ bits of space.
Notice that the space complexity is optimal up to constant factors, placing the derandomization precisely in $\mathbf{L}$, as long as the alphabet bit-length $s=O(\log w)$. Also notice that this implies our algorithm is in time $\operatorname{\mathrm{poly}}(w)$. For derandomizing regular ROBPs within $\mathbf{L}$, the only prior feasible methods are Nisan-Zuckerman generator [NZ96] and Armoni’s generator [Arm98]. Our result significantly improves the length of the regular ROBPs that can be derandomized in $\mathbf{L}$ from $\operatorname{\mathrm{poly}}\log w$ to $2^{O(\log^{1/2}w)}$ and also improves the precision from $2^{-\log^{0.99}w}$ to arbitrary $1/\operatorname{\mathrm{poly}}w$.
If compared to the Saks-Zhou style derandomizations of [AKM+20, CHL+23, CL24], then our algorithm not only attains optimal space complexity $O(\log w)$, but also has a substantial improvement for time complexity from super-polynomial to standard polynomial, as long as $n=2^{O(\log^{1/2}w)}$. Note that derandomizations of [AKM+20, CHL+23, CL24] are not in polynomial time even if $n=\operatorname{\mathrm{poly}}(\log w)$ and $\varepsilon=O(1)$.
@@ 1.5 Technical Overview
WPRGs for standard-order ROBPs
We start by reviewing the Richardson iteration based error reduction [AKM+20, CDR+21, PV21] that attains low-error WPRGs from large-error PRGs.
The Richardson iteration based error reduction can be viewed as a procedure that produces high-precision approximations for iterated matrices multiplications, given low-precision approximations. Let $\{A_{i}\}_{i=1}^{n}\subseteq\mathbb{R}^{w\times w}$ be transition matrices, and $\{B_{i,j}\}_{0\leq j<i\leq n}\subseteq\mathbb{R}^{w\times w}$ be approximations satisfying $\|B_{i,j}-A_{i}\cdots A_{j+1}\|\leq\varepsilon_{0}/(2(n+1))$. For any integer $k$, a standard method, e.g. [CDR+21], indicates there is a weighted sum $P=\sum_{i=1}^{K}\sigma_{i}B_{n_{i,1},n_{i,2}}\cdots B_{n_{i,k-1},n_{i,k}}$ of $K=n^{O(k)}$ terms such that
$\left\|A_{n}\cdots A_{1}-P\right\|\leq\varepsilon_{0}^{k/2}(n+1),$
where $\sigma_{i}\in\{-1,0,1\}$ and $\forall i,$ the sequence $n_{i,j},j\in[k]$ is an increasing sequence of indices in $[n]$. [^4] See Appendix B for details. To apply this to an ROBP $f$, one can instantiate the matrices as
  • $A_{i}$: the stochastic matrix of $f$ from layer $i-1$ to $i$.
  • $B_{i,j}:=\mathbb{E}_{x\sim\{0,1\}^{s}}\left[f^{j\to i}(\mathsf{PRG}(x))\right]$, where $f^{j\to i}$ is the transition of $f$ from layer $j$ to layer $i$, and $\mathsf{PRG}$ is an $\frac{\varepsilon_{0}}{2(n+1)}$-PRG with seed length $s$.
This immediately yields a high-precision approximation of the expectation of the ROBP:
$\left|\mathbb{E}_{U}[f(U)]-\sum_{i=1}^{K}\sigma_{i}\cdot\mathbb{E}_{x_{1},\dots,x_{k}\sim\{0,1\}^{s}}f\left(\mathsf{PRG}_{0\to n_{i,1}}(x_{1}),\dots,\mathsf{PRG}_{n_{i,k-1}\to n_{i,k}}(x_{k})\right)\right|\leq\varepsilon_{0}^{k/2}(n+1).$
Notice that this also immediately gives a WPRG for $f$ with error $\varepsilon_{0}^{k}(n+1)$ by definition. However the $k$ independent PRG callings cost too much randomness. [CDR+21] further reduces the randomness by using an INW generator to generate the seeds for the $k$ independent PRG callings. But since INW does not have optimal seed length, this takes $O(\log k\log\frac{kw}{\varepsilon}+s)$ randomness which has an extra $O(\log k)$ factor in the main term. Using the sampler technique of [Hoz21], one can avoid this factor, but only when the starting error $\varepsilon_{0}$ is already $1/\operatorname{\mathrm{poly}}w$. If the large error PRG has $\varepsilon_{0}\gg 1/\operatorname{\mathrm{poly}}w$, then this does not work as desired. For example, if the large-error PRG is the NZ PRG, then it is unclear how to use this to derive an optimal seed length for $\varepsilon=1/\operatorname{\mathrm{poly}}w$.
To address the problem, we introduce a new strategy. For each $i\in[K]$, one can view
$f_{i}(y_{1},\dots,y_{k}):=f\left(\mathsf{PRG}_{0\to n_{i,1}}(y_{1}),\mathsf{PRG}_{n_{i,1}\to n_{i,2}}(y_{2}),\dots,\mathsf{PRG}_{n_{i,k-1}\to n_{i,k}}(y_{k})\right),$
as a new ROBP with a shorter length $k$ and a large alphabet, while the width is still $w$. To fool $f$, one only need to fool these shorter ROBPs. Based on this observation, our high-level idea is to iteratively apply the above procedure to reduce the ROBPs to even shorter ones until the length become a constant such that the reduced ROBPs can be fooled trivially.
We use the notion weighted pseudorandom reduction to describe the construction in details.
@@
Let $\mathcal{B}_{n_{0},s_{0},w}$ be the class of length-$n_{0}$, width-$w$ ROBPs over alphabet $\Sigma_{0}=\{0,1\}^{s_{0}}$. Let $\mathcal{B}_{n_{1},s_{1},w}$ be the class of length-$n_{1}$, width-$w$ ROBPs over alphabet $\Sigma_{1}=\{0,1\}^{s_{1}}$. A $(d,K,\varepsilon)$-reduction from $\mathcal{B}_{n_{0},s_{0},w}$ to $\mathcal{B}_{n_{1},s_{1},w}$ is a tuple $(\mathsf{R},\sigma)$, in which $\mathsf{R}:\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\times\{0,1\}^{d}\rightarrow\left(\{0,1\}^{s_{0}}\right)^{n_{0}}$ and $\sigma:\{0,1\}^{d}\rightarrow\mathbb{R}$. Furthermore, for every $f\in\mathcal{B}_{n_{0},s_{0},w}$, we have:
@@
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d}}\sum_{i\in\{0,1\}^{d}}\sigma(i)\mathbb{E}_{U^{\prime}}[f(\mathsf{R}(U^{\prime},i))]\right|\leq\varepsilon,$
$\forall i,|\sigma(i)|\leq K,$
$\forall i,f(\mathsf{R}(\cdot,i))\in\mathcal{B}_{n_{1},s_{1},w}.$
We call $\mathsf{R}$ the reduction function and $\sigma$ the weight function.
Let the original length-$n$, width-$w$ ROBP be $f$ over alphabet $\Sigma$. For simplicity, now let the target error $\varepsilon=1/\operatorname{\mathrm{poly}}w$. We plan to construct a sequence of length reductions as the following. Let $n_{0}=n$. The $i$-th length reduction, which is instantiated by the Richardson iteration based error reduction with a base $\frac{\varepsilon_{i}}{2(n_{i-1}+1)}$-PRG, is a $(d_{i},K_{i},\varepsilon)$-reduction from $\mathcal{B}_{n_{i-1},s_{i-1},w}$ to $\mathcal{B}_{n_{i},s_{i},w}$, where we take
$n_{i}=O\left(\frac{\log(n_{i-1}/\varepsilon)}{\log(1/\varepsilon_{i})}\right).$
This may introduce error
$\varepsilon_{i}^{n_{i}/2}(n_{i-1}+1)\leq\varepsilon,$
as it is instantiated by the Richardson iteration based error reduction. Notice that as $i$ goes up, one can decrease $\varepsilon_{i}$ such that $n_{i}$ can be smaller and smaller. One may doubt that a smaller $\varepsilon_{i}$ may require a more precise base PRG, which may require a longer seed. However, the seed length depends on $n_{i-1},w,\varepsilon_{i}$. As $n_{i-1}$ is already decreased by the previous reduction, one can pick a smaller $\varepsilon_{i}$ without increasing the seed length. Another issue is that the alphabet is accumulating quickly as the seed length of the base PRG will become the alphabet bit-length for the next reduction. But we shall see later that one can add an alpahbet reduction after each length reduction to retain a small alphabet.
More concretely, the $l$ length reductions mentioned above are instantiated by the Richardson iteration based error reduction with Armoni’s generator [Arm98, KNW08] as base PRGs. For a length-$n$, width-$w$ ROBP over $\Sigma=\{0,1\}^{s}$, we first choose $\varepsilon_{0}=2^{-\sqrt{\log w\log n}}$ and apply a $(d_{0}=\log K_{0},K_{0},\varepsilon)$-reduction from $\mathcal{B}_{n,s,w}$ to $\mathcal{B}_{n_{1},s^{\prime}_{1},w}$, in which $n_{1}=O\left(\sqrt{\frac{\log w}{\log n}}\right)$, $K_{0}=2^{O\left(\sqrt{\log n\log w}\right)}$ and $s^{\prime}_{1}=O\left(s+\frac{\log n\log(nw/\varepsilon_{0})}{\log\log w-\log\log(n/\varepsilon_{0})}\right)$. For the subsequent $l-1$ reductions i.e. for $i=1,2,\ldots,l$, we setup a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$ reduction from $\mathcal{B}_{n_{i},s_{i},w}$ to $\mathcal{B}_{n_{i+1},s^{\prime}_{i+1},w}$, where we assume $s_{i}=c\log w$ with $c$ being a universal constant. We choose $\varepsilon_{i}=2^{-\log w/n_{i}^{1/3}}$, and consequently $n_{i+1}=n_{i}^{1/3}$, $K_{i}=O\left(2^{n_{i}^{1/3}\log n_{i}}\right)=o\left(\log^{1/2}w\right)$ and $s^{\prime}_{i+1}=O(s_{i}+\log w)$. We iterate for $l$ steps until we reach $n_{l}=O(1)$. One can see that $l\leq O(\log\log\log w)$.
As mentioned before, for every $i$, after the $i$th length reduction, we apply an alphabet reduction to retain a small alphabet. Each alphabet reduction is realized by a one-level NZ generator. Assume we have a ROBP $f^{\prime}$ with length $m=n_{i}$ and alphabet bit-length $s^{\prime}$. The one-level NZ generator which, given a source $X$ of $n=O(s^{\prime}+\log w)$ bits and a sequence of independent seeds $Y_{1},\ldots,Y_{m}$, each of $d=O(\log(n/\varepsilon))$ bits, outputs $(\mathrm{E\textsc{xt}}(X,Y_{1}),\mathrm{E\textsc{xt}}(X,Y_{2}),\ldots,\mathrm{E\textsc{xt}}(X,Y_{m}))$ where each $\mathrm{E\textsc{xt}}(X,Y_{i})$ is of binary length $s^{\prime}$. Now fixing $X=x$, the function $f^{\prime}\left(\mathrm{E\textsc{xt}}(x,y_{1}),\mathrm{E\textsc{xt}}(x,y_{2}),\ldots,\mathrm{E\textsc{xt}}(x,y_{m})\right)$ can be viewed as an ROBP on input $(y_{1},\ldots,y_{m})$ with alphabet bit-length $d=O\left(\log\frac{s^{\prime}+\log w}{\varepsilon}\right)$. One can check $d\leq c\log w$ as $\varepsilon=1/\operatorname{\mathrm{poly}}w$, where $c$ is a universal constant.
The seed length of our WPRG is the some of the randomness of all reductions, together with the true randomness used to fool the final constant length ROBPs. For length reductions the seed length is $\sum_{i=0}^{l}d_{i}\leq O\left(\log(nw)\right)$. The first alphabet reduction contributes $O\left(s+\frac{\log n\log(nw)}{\max\left\{1,\log\log w-\log\log n\right\}}\right)$. The remaining alphabet reductions contribute $O(l\cdot\log w)$. The true randomness for final constant length ROBPs is $O(\log w)$ since the alphabet bit-length is $O(\log w)$.
Finally for even smaller errors ($\varepsilon\ll 1/\operatorname{\mathrm{poly}}(w)$), we apply a variant of Hoza’s sampler [Hoz21] technique, which can also be viewed as an extra level of weighted pseudorandom reduction.
We start by reviewing the recent WPRG given by [CHL+23] in the view of weighted pseudorandom reductions. Their WPRG combines the INW generator and a new matrix iteration method. To fool a unbounded-width permutation ROBP of length $n$ and alphabet $\{0,1\}^{s}$ within error $\varepsilon$, they first prepare a base INW generator $\mathsf{PRG}:\{0,1\}^{d}\to\left(\{0,1\}^{s}\right)^{n}$ with moderate sv-error [^5] See Definition 4.5. $\tau=2^{O(\log\log n\sqrt{\log(1/\varepsilon)})}$. Then they construct the $\varepsilon$-WPRG $(G,\sigma)$ in the following form:
@@
$\sigma(j)=\sigma_{j}\cdot K,$
@@
where $j\in[K]$, $K=2^{O(m)}$, $m=O\left(\log n\cdot\frac{\log(1/\varepsilon)}{\log(1/\tau)}\right)$, for each $j$ the sequence $n_{j,k},k\in[m]$ of indices being such that $0\leq n_{j,1}<n_{j,2}<\ldots<n_{j,m}\leq n$, and $\sigma_{j}\in\{-1,0,1\}$. The seed length of $\mathsf{PRG}$ is $d=s+O\left(\log n\cdot(\log\log n+\log(1/\tau))\right)$. Note that for each $j\in[K]$, one can view $f(G(j,x_{1},\ldots,x_{m}))$ as a new permutation ROBP on input $x_{1},x_{2},\ldots,x_{m}$. [CHL+23] applies an $\varepsilon/K$-INW generator to fool such reduced permutation ROBPs, which cost $d+O\left(\log m\cdot(\log\log m+\log(K/\varepsilon)\right)$ bits of randomness. This step introduces some double-logarithmic factors.
To address this problem, we introduce the following new strategy based on weighted pseudorandom reductions. Let $\mathcal{P}_{n,s}$ denote the class of length-$n$ unbounded-width permutation ROBPs over alphabet $\{0,1\}^{s}$. Then $(G,\sigma)$ can be viewed as a $(\log K,K,\varepsilon)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{m,d}$. Let our target error be $\varepsilon$. We setup $l$ reductions as the following. The $i$-th reduction is a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$. For the first reduction, we set $\tau_{1}=O\left(\sqrt{\log(1/\varepsilon)}\right)$ to obtain a $(\log K_{1},K_{1},\varepsilon/2)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{n_{1},s_{1}}$. One can see $n_{1}=O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)$, $s_{1}=s+O\left(\log n\cdot\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)\right)$ and $K_{1}=\exp\left(O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)\right)$. For the subsequent $l-1$ reductions, we set $\varepsilon^{\prime}=(\varepsilon/(2K_{1}))^{2}$ and $\tau_{i}=2^{\frac{C\log(1/\varepsilon^{\prime})}{\log^{2}n_{i-1}}}$ to obtain a$(\log K_{i},K_{i},\varepsilon^{\prime})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$, where $C>0$ is a universal constant. One can deduce that $n_{i}=O\left(\log n_{i-1}\frac{\log(1/\varepsilon^{\prime})}{\log(1/\tau_{i})}\right)=\log^{3}n_{i-1}$, $s_{i}=s_{i-1}+O(\log n_{i-1}\cdot(\log\log n_{i-1}+\log(1/\tau_{i})))=s_{i-1}+O\left(\frac{\log(1/\varepsilon^{\prime})}{\log n_{i-1}}\right)$ and $K_{i}=2^{O(\log^{3}n_{i})}$. We do the iteration until $n_{l}=O(1)$. So $l=o(\log\log n)$.